OpenStudy (anonymous):

integral from 0 to two ∫² 2t÷(t-3)²

OpenStudy (anonymous):

yest i did but what happens to the 2t

OpenStudy (anonymous):

i dont understand how to integrate

OpenStudy (anonymous):

its easy, just takes some practice

OpenStudy (anonymous):

u=t-3 du=dt u+3=t 2u+6=2t when t=0 u=-3 when t=2 u=-1 \[\int\limits_{-3}^{-1}2u+6/u^2du\] \[\int\limits_{-3}^{-1}2/u+6/u^2du=2\ln|u|-6/u|[-1,-3]\] 2ln(1)+(-6/-1)-(2ln(3)+(-6/-3)= 0+6-2ln(3)-2 4-ln(9)=number but thats the purest answer

OpenStudy (anonymous):

[-3,-1] not the opposite

OpenStudy (anonymous):

\[\int\limits_0^2 \frac{2t}{(t-3)^2} \, dt=4-\text{Log}[9] \]